OpenStudy (danbaba):

0=2ln(x/4) + 1 I don't understand how to solve this problem. Thank you!

OpenStudy (danbaba):

Wolfram is telling me \[x=\frac{ 4 }{ \sqrt(e) }\] But I don't understand HOW to find this.

OpenStudy (sooobored):

ok, what you want to do is isolate x onto one side or solve for x

OpenStudy (sooobored):

starting with \[2 \ln(\frac{x}{4})+1=0\]

OpenStudy (sooobored):

first we want to remove the 1 from the left side this can be done by subtracting 1 to both sides, since 1-1 =0 so it will look something like \[2 \ln(\frac{x}{4})+1-1=0-1\] \[2 \ln(\frac{x}{4})=-1\]

OpenStudy (danbaba):

100% following. (I pretty much was stuck at: \[\ln(x) = -\frac{ 1 }{ 2 } + \ln(4)\] Which I think is correct that far.

OpenStudy (sooobored):

er, i wouldnt do it like that

OpenStudy (danbaba):

Oh ok, I figured I was making this more difficult. Haha, please continue I won't interrupt.

OpenStudy (sooobored):

i would then divide by 2 from both sides\[\frac{2}{2} \ln(\frac{x}{4}) = \frac{-1}{2}\] so \[ \ln(\frac{x}{4}) = \frac{-1}{2}\]

OpenStudy (sooobored):

then take both sides to be the power of the exponential function since e^(lnx) = x

OpenStudy (sooobored):

\[e^{\ln{\frac{x}{4}}}=e^\frac{-1}{2}\] \[\frac{x}{4}=e^\frac{-1}{2}\]

OpenStudy (danbaba):

omg, yep... I see it now thanks to you. I dunno why I was thinking x/4 couldn't do that...

OpenStudy (sooobored):

now \[x^{\frac{-1}{2}} = \frac{1}{\sqrt{x}}\]

OpenStudy (sooobored):

then we just multiply 4 to both sides and you should get the answer above

OpenStudy (danbaba):

Thanks so much.

OpenStudy (sooobored):

yup no prob